A merry go round is spinning in a circle at a constant speed. A child has a centripetal acceleration of 10 m/s^2 and a tangential velocity of 5m/s, what is the radius of the merry go round?
1
Expert's answer
2020-12-15T11:46:47-0500

The centripetal acceleration can be expressed as


"a=\frac{v^2}{R},\rightarrow R=\frac{v^2}{a}=2.5\text{ m}."